What is the equation of the line that contains the points with (x, y) coordinates (–3, 7) and (5, –1)?
Answer by lwsshak3(11628)

What is the equation of the line that contains the points with (x, y) coordinates (–3, 7) and (5, –1)
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
For given equation:
m=∆y/∆x=7-(-1)/5-(-3)=8/8=1
y=(1)x+b
solving for b using (x,y) coordinates of one of given points (5,-1)
-1=5+b
b=-6
equation:
y=x-6